# Self-organisation - the underlying principle and a general model

### v2

@@ Abstract
Recent observations of coordinated self-organisation (SO) of stress and structure in granular systems provide insight into the fundamental principle underlying this phenomenon. It is first argued here that SO emerges when a minute subset of configurations are significantly more stable than the rest and therefore survive the noise in the system much longer to be observed. This principle goes deeper than recently proposed energy considerations. Guided by this principle, a statistical mechanics model is formulated then for SO in these systems and its extension to three dimensions is outlined. The principle holds beyond granular systems and the model is extended next to describe emergence of SO in more general systems. The application of the model is illustrated for the specific example of laning. Parallels of the modelling approach to traditional statistical mechanics provide useful insight that should assist in modelling SO in other out-of-equilibrium systems.
@@ I. Introduction
Most studies of SO have been usually focused on specific contexts and in many case are descriptive. Attempts at finding a general principle underpinning SO are few and they are mainly based on energy considerations. For example, it has been proposed that SO emerges in steady states by selection of microstates that minimise energy-dissipation in driven systems Cyetal21 ; Keetal23 . These models propose that the system random-walks in the multi-dimensional microstates space and those special microstates are selected as the system self-organises. However, this does not explain what is exactly the physical mechanism that makes the random walk park on the microstates of low energy dissipation. Moreover, this idea excludes emergence of SO in out-of-equilibrium a-thermal systems, where energy plays a negligible role. Observations of SO in these systems suggest that a more general principle is in play, which applies to all systems, however driven. Statement of the principle is insufficient - it should also point the way toward a unified modelling approach that leads to prediction of the self-organised state and some relevant properties.
The first aim here is to propose a more general principle that applies beyond energy considerations. The second aim is to use the principle to formulate a framework of modelling SO in general. For didactic reasons, the principle is illustrated initially in the context of a specific example - driven two-dimensional (2D) granular systems. Then, a statistical-mechanics-based model is formulated for these systems and its extension to three-dimensional systems is outlined. A discussion how the principle applies more generally follows and it pave the way for a corresponding statistical mechanics formalism of SO in general. The application of the general formulation to a simple example of SO, where energy plays no role, is detailed in the supplemental material. In the concluding discussion, similarities to, and differences from, thermal systems in equilibrium are presented. These substantiate the generality of the principle and, more importantly, the parallels provide insight into statistical mechanics modelling of SO in other contexts. Future extensions are also proposed.
II. SO signatures in 2D granular systems
In steady states of out-of-equilibrium thermal systems, the invested driving energy is balanced against internal dissipation. The noise in these processes is generated both by the dynamics and, depending on magnitude, also by thermal fluctuations. To illustrate the thesis here, I begin with clarifying the principle in the context of the specific example - the emergence of SO in quasi-static dynamics of driven 2D granular systems. In recent experiments and simulations, the driving methods were simple shear, pure shear, and both isotropic and anisotropic compression. The term quasi-static means that any dynamic process is slower than the response time of stress equilibration, which also means that stopping at any moment, the resulting structural configuration of the assembly of particles is in mechanical equilibrium. It is these configurations that display signatures of SO. Almost all these signatures were seen by studying one structural element - the cells. These are the smallest possible voids in the assembly, which are surrounded by particles in contact. The first signature is that distributions of several structural quantities collapse onto a master forms, once simply scaled by their means. Such a collapse was observed for many (up to 120!) different systems and different packing protocols MaBl14 . Another intriguing signature involves the evolution of cell orders, where an order is defined as the number of particles in contact around a cell. Cell orders change when contacts are made and broken and these processes resemble chemical reactions. It has been found both numerically and experimentally that the dynamics evolve toward steady states that satisfy detailed balance Waetal20 ; Suetal25 . Another observation, providing a clue to the origin of the SO, was of a competition between the stability of cells and their configurational entropy MaBl17 ; Suetal20 . The smoking gun, however, has been observations that, in a number of systems and protocols, there are strong correlations across the systems between the local principal axes of cell shapes and cell stresses MaBl17 ; MaBl21 ; Jietal25 . This strong local alignment demonstrates clearly that the local structure and local stress self-organise cooperatively, a feature that has to be included in any potential model of SO.
These observations offer insight into the principle that underlies the SO in these 2D granular dynamics. The driving and the boundary conditions generate an internal stress field that fluctuates spatially and temporally. If a cell cannot support the local stress it disintegrates, namely, at least one of the contacts of the particles surrounding it breaks. A system-wide configuration consists of a snapshot of all its cell configurations. The SO phenomenon arises when the occurrence probabilities of a minute subset of cell shapes, orientations, and stresses configurations dominates the observations. The reason for this dominance is that these configurations of cell shapes and stresses are much more stable than the rest and therefore survive much longer. All other configurations are not observed in simulations and experiments because their lifetimes are too short. In the granular systems, the cell shape and cell stress are locally coupled. When unstable cells either rearrange or disintegrate, the stress around them also resets. Thus, local stresses persist for longer around more stable cells and stability is correlated with the large principal stress aligning with the cell’s major axis. This interpretation is consistent with all the numerical and experimental observations and, in particular, it is almost surely the reason for the strong system-wide stress-shape orientational correlations.
@@ III. SO statistical mechanics in granular systems
With occurrence probabilities of configurations playing a key role in SO emergence in these dynamics, a natural theoretical model for it is statistical mechanics. To formulate it, several definitions need to be introduced. A cell surrounded by $k$ particles in (force-carrying) contact is a cell of order $k$, or $k$-cell. The cell order distribution across the system is discrete, with $Q_{k}$ (k=3,4,,Kk=3,4,...,K) the fractions of $k$-cells out of all $N_{c}$ cells. Cell stability depends on two ’attributes’: its shape and its stress. To quantify the former, cells are approximated as ellipses Jietal25 . A cell ellipse has three degrees of freedom (DOFs): its major and minor axes lengths, $a_{c}$ and $b_{c}$, respectively, and the orientation of the major axis relative to a fixed frame of reference, $\theta_{c}$. Cell stresses are defined, using an appropriately weighted sum over the stresses of the particles surrounding it Bletal25 , as detailed in the supplemental material SM . The cell stress also has three degrees of freedom: the large and small principal stresses, $\sigma_{c1}$ and $\sigma_{c2}$, respectively, and the orientation of the large principal stress relative to the same fixed frame of reference, $\phi_{c}$.
At the heart of any statistical mechanical formalism is a partition function, which accounts for the occurrence probability of each system-wide configuration. In this approach, cells are regarded as quasi-particles, a configuration’s occurrence probability should increase with its stability, which would also increase its survivability. A cell’s stability depends on both its shape and its stress. The smaller the difference between the its ellipse axes lengths and its principal stresses the stabler it is. Cell stability plays a central role and stability measures need to be quantified. A standard measure of stability is VMi13
@@
Any model for the shape stability can be used and, for illustration purposes, a similar form is assigned here,
@@
Next, a survivability function is required, whose value determines the configurations occurrence probabilities. It should include the stability measures, as well as any other constraints on the system. It’s main aim is to quantify the stability, and thence survivability, of system-wide configurations. The observations of function collapses and correlations between stress and shapes further suggest coupling between $\theta_{c}$ and $\phi_{c}$ at the cell level, which the survivability function should include. Additionally, it has been found Jietal25 that the most probable values of the stability measures are finite, which is another piece of information that needs to be accommodated in the survivability function. In principle, it could also include correlations between the DOFs of nearest-neighbour cells, although no such correlations have been observed in the experimental and numerical systems Waetal20 ; Suetal25 . The survivability function should decreases as cell stability increases, which would make it analogous to the free energy in thermal systems. Finally, the partition function should also be grand-canonical because the number of cells, $N_{c}$, fluctuates. All these specifications are satisfied by
@@
In this expression, $\mu$, $J_{1}$, and $J_{2}$ are Lagrange multipliers that reflect the effect of the dynamics-generated noise on the number of cells, the cell shape stability, and the cell stress stability, respectively. The last term represents the cell-cell interaction, with $\psi_{c^{\prime}c}=\left(\theta_{c}-\theta_{c^{\prime}}\right)^{2}$ the difference between the orientations of neighboring cells $c$ and $c^{\prime}$. More interaction terms between other DOFs of neighbouring cells can be added to $F_{c}$, if necessary.
As the DOFs are continuous, it makes sense to convert the summation in (3) into integrals of ‘densities of states’ over the phase space of DOFs. Denoting the cell DOFs by
$\left\{X\right\}=\left\{a,b,\theta,\sigma_{1},\sigma_{2},\phi\right\}\ ,$
the densities of states provide the occurrence probabilities of the DOFs and/or the cells stability measures. For the 2D granular systems, some of these distributions are known from modelling and observations. Specifically, $\theta_{c}$ is isotropic on the cell level and $\phi_{c}$ is strongly correlated with it Jietal25 . Thus, these two variable can be replaced by one variable, $\Delta\equiv\theta-\phi$. This variable is normally distributed almost identically around zero for all $k$, $P(\Delta)\sim e^{-\Delta^{2}/\left(2\delta^{2}\right)}$. In the absence of other information, the distributions of the stability measures $h$ and $f$ can be obtained from the distributions of the DOFs on which they depend. However, it has been shown that, in these systems, the conditional distributions of $h_{c}$ given $k$, $P_{h}(h|k)$, when scaled by the mean for each cell order, $\bar{h}(k)$, collapse onto an identical Weibull form, $W\left[\hat{h}=h/\bar{h}(k)\right]$, that is independent of $k$ Jietal25 . Observations suggest that, at least in two shear experiments, there are no cell-cell spatial correlations Waetal20 ; Suetal25 . Namely, $J_{c^{\prime}c}=0$ and this becomes an ‘independent cells model’. To demonstrate the use of the formalism, assume for simplicity that the conditional densities of states of $f(k)$ also collapse onto a master form under a similar scaling. An explicit calculation of the partition function, detailed in the supplemental material SM , yields
@@
with the angular brackets denoting averages over all the DOFs. This expression can be used to obtain expectation values of quantities, which depend on the DOFs and the stability measures, as functions of the noise parameters and coupling constants in the survivability function,
@@
E.g., analogously to traditional statistical mechanics, one can readily derive
$\langle f\rangle=\sum\limits_{k=3}^{K}Q_{k}\bar{f}(k)=-\frac{\partial\ln{Z}}{\partial J_{1}}$ (7)
$\langle h\rangle=\sum\limits_{k=3}^{K}Q_{k}\bar{h}(k)=-\frac{\partial\ln{Z}}{\partial J_{2}}\ .$ (8)
The extension of this formalism to three-dimensional (3D) granular SO is straightforward in principle. As in 2D, cells are the smallest polyhedral voids surrounded by particles in contact. They can be approximated as ellipsoids and their stresses defined as weighted sum of the stresses of their surrounding particles. This would give 10 DOFs per cell (three axes and two angles for each of the shape and stress ellipsoids). Similar ellipsoid and stress stability measures can be defined and $F_{c}$ can be a straightforward 3D version of (4). This, in principle, provides the partition function. However, its evaluation may be more difficult than in 3D, especially because of the more complicated cells classification into families. Cells are polygons in 2D and polyhedra in 3D and, while a polygon can be classified only by its order, classification of polyhedra requires more parameters: the number of its faces and the order of each polygonal face. Specifically, a family of cell polyhedra is classified by its $n_{p}$ vertices (contacts), $k_{3}$ triangular faces, $k_{4}$ rectangular faces, and so on up to the highest-order faces, $k_{t}$. Members of the same family have an identical series of such numbers. Nevertheless, this complication is a quantitative, rather than qualitative. It just means that the product over $k$ in (3) needs to be replaced by products over all the cell parameters, but the integrand remains the same. This extension is under investigation currently, but it is downstream from the aim here and will be presented elsewhere.
@@ IV. The principle and its generalisation
Emergence of SO in the above example is governed by a clear principle. Cells that are better aligned with the local principal stress direction are more stable and survive longer the fluctuating local stress. In turn, stress fluctuations result from cell rearrangements and disintegration. Thus, evolving cooperatively lengthens the lifetime of both cell stress and structure configurations. The longer local configurations survive the more frequently observed are system-wide states comprising these local configurations. This interpretation is supported by numerical and experimental observations in disordered disc assemblies. It was found that the mean area of $k$-cells, when normalised by the area of the regular $k$-polygon of the same edge length, first decreases with $k$ and then it increases again MaBl17 ; Suetal20 . The initial decrease is the result of an increasing number of elongated cell configurations, which are stable at low cell orders but become increasingly unstable as $k$ increases. This means that stress and structure are coupled and cell stability limits configurational entropy.
This principle can be extended beyond granular systems. Driving a disordered system generates noise even if thermal fluctuations are negligible. Such systems are also constrained by the environment through boundary conditions and interactions with other dynamic systems, as well as internally by the nature of interactions between their individual components. SO emerges when only a minute subset of system configurations, its microstates, survive the noise sufficiently long to dominate experimental observations. In contrast to these, the bulk of system-wide configurations disintegrate faster than observational timescales and therefore are not typical. As it is the driving that determines the noise and it is the noise that determines which configurations survive, different driving methods may give rise to SOs of different characteristics.
V. Extension of the modelling framework
The extension to more general modelling of SO is as follows. (1) define quasi-particles (e.g., the cells); (2) identify the quasi-particles’ DOFs (the cell ellipse axes lengths and orientations, principal stresses and orientation); (3) identify quasi-particle stability measures that depend on the DOFs and whose values decrease with increasing quasi-particle stability (the cell stresses and shapes); (4) construct a survivability function that depends on all the quasi-particles’ stability measures and decreases monotonically as their stability increases ($F_{c}$ in eq. (4)). The value of the survivability function at each system-wide configuration is determined by the configuration’s stability, which, in turn, determines the configuration’s occurrence probability and therefore its lifetime.

### v3

@@ Abstract
It is proposed that self-organisation (SO) in driven systems is governed by a general principle: it emerges when a minute subset of system configurations are exceptionally stable and long-lived to survive the noise generated by the driving. Guided by this principle, a statistical mechanics-like model is formulated for general SO and its application is illustrated for two example systems: self-organised steady states of quasi-statically driven granular systems in two dimensions and crowd laning, for which illustrative explicit results are derived. In this formalism, maximising a survivability function of the exceptionally few stable configurations is the equivalent of minimising the free energy in traditional statistical mechanics. The parallels with equilibrium statistical mechanics provide useful insight, which should assist in modelling SO in general out-of-equilibrium systems. Similarities and differences between SO in passive and biological systems are also pointed out, suggesting potential extension of this approach in this direction, albeit to very simple systems.
@@ I. Introduction
Most studies of SO have been usually focused on specific contexts and in many case are descriptive. Attempts at finding a general principle underpinning SO are few and they are mainly based on energy considerations. For example, it has been proposed that SO emerges in steady states by selection of microstates that minimise energy-dissipation in driven systems Cyetal21 ; Keetal23 . These models propose that noise ’rattles’ the system and, as it is bumped from microstate to microstate, it gets trapped in special microstates where the noise is very low. Yet, the physical mechanism for the selective low rattling in those microstates is unclear. A physical mechanism is suggested in Keetal23 , based on inability to escape energy potential wells. However, it is unclear if this explains other energy-dissipating systems and how to extend it to SO in athermal systems, where thermal noise is negligible compared with noise from other sources, especially when energy-based considerations plays little to no role in determining the self-organised state. Observations of SO in the latter systems suggest that a more general principle is in play, which applies to all systems, however driven. Motivated by this idea, the main aim here is to state the fundamental principle that governs SO phenomena in general and, based on it, propose a formalism to determine their characteristics. The approach is analogous to statistical mechanics, a method that has been applied in recent decades well beyond thermal equilibrium, e.g. EdOa89a ; EdOa89b ; Ga99 ; Vietal99 ; Ya09 ; Taetal24 . Like traditional statistical mechanics, the formalism can predict the characteristics of the long-term self-organised steady states in general - it does not model the emergence SO from specific driving dynamics. A significant advantage of this approach is that the parallels between the proposed formalism and traditional statistical mechanics offer insight into better modelling of SO in non-equilibrium systems.
For didactic reasons, the principle is first illustrated in the context of a specific example - driven granular systems, and its generalisation is discussed. For the same reason, the approach is also formulated initially in the context of this system. While the formalism is general, to demonstrate its predictive power in specific systems, some assumptions must be made and these are stated, under which expectation values and an equivalent of an equation of state are derived explicitly. The general method is then formulated and discussed. This is followed by another example that illustrates the use of the formalism - a simple model of crowd laning SO. Both examples make clear the applicability of the principle and the method to SO that is not determined by energy considerations. A discussion follows of the applications of the formalism, in principle, to biological systems. The paper concludes with a summary, a discussion of similarities and differences between this method and equilibrium thermal systems, and possible future work.
II. SO in 2D granular systems and the principle
Recent observations of SO in experiments and simulations of 2D granular systems provide the key to identify the principle governing this phenomenon MaBl14 ; Waetal20 ; Suetal25 ; MaBl17 ; Suetal20 ; MaBl21 ; Jietal25 . Driven quasi-statically out of equilibrium by various methods and for a wide range of parameters, it has been found that: many measurable quantities appeared to collapse to master curves MaBl14 , structural transitions converge to detailed-balance-like steady states Waetal20 ; Suetal25 , and local stress and structural configurations self-organise cooperatively MaBl17 ; Suetal20 ; MaBl21 ; Jietal25 . All these signatures involved the cells, which are the smallest possible voids surrounded by particles in contact. The latter is very significant; when cells were approximated as ellipses, their orientations were strongly correlated with the local principal stress axes. As cell orientations were isotropic and independent of the boundary conditions, this was strong evidence of a cooperative stress-structure SO. The physics behind this correlation is not difficult to understand. The driving generates noise, causing the internal stress field to fluctuate in space and time. A cell that cannot support the local stress disintegrates, i.e., at least one of the contacts of the particles surrounding it breaks. This causes the stress around it to reconfigure and stabilise when a new cell is formed. The stablest configuration corresponds to the cell’s ellipse aligning with the largest principal local stress (think of resistance to pressing on an egg along its long or short axes). Being stable, these cell shape-stress configurations are sufficiently long-lived to be observed consistently, i.e., the system self-organises into them. This sheds light on a general principle that governs SO - a minute subset of configurations dominates observations because of their higher stability and longer survivability. Extending this interpretation to a general principle is discussed below.
@@ III. SO statistical mechanics in granular systems
As occurrence probabilities of long-term configurations are at the core of SO, a natural modelling approach is statistical mechanics - a method that derives system-wide properties on the basis of such probabilities. For the granular systems, cells are chosen as quasi-particles because all the aforementioned signatures involved these structural descriptors. Being defined by the contacts of particles around them, cells have the advantage of sensitivity to the local forces, which are transmitted via the contacts. It is this sensitivity that gives rise to the strong correlations between cell shapes and local stresses Jietal25 . Descriptions based on Voronoi tessellations, for example, smear this correlation because they are based on particle proximity rather than contacts. The actual grains are also less useful as quasi-particles because their shapes are fixed and their characteristics are not as sensitive to local stress fluctuations. A cell surrounded by $k$ particles in contact is a cell of order $k$, or $k$-cell. The cell order distribution across the system is discrete, with $Q_{k}$ (k=3,4,,Kk=3,4,...,K) the fractions of $k$-cells out of all $N_{c}$ cells. Cell stability depends on two ’attributes’: its shape, approximated as an ellipse Jietal25 , and its stress. A cell ellipse has three degrees of freedom (DOFs): the major and minor axes lengths, $a_{c}$ and $b_{c}$, respectively, and the orientation of the major axis relative to a fixed frame of reference, $\theta_{c}$. Cell stresses are defined, using an appropriately weighted sum over the stresses of the particles surrounding it Bletal25 , as detailed in the supplemental material SM . In 2D, the cell stress also has three DOFs: the large and small principal stresses, $\sigma_{c1}$ and $\sigma_{c2}$, respectively, and the orientation of the large one relative to the same fixed frame of reference, $\phi_{c}$.
At the heart of the statistical mechanics formalism is a partition function, which accounts for the occurrence probability of each system-wide configuration. In the proposed approach, the more stable the cell’s shape and stress the higher is its occurrence probability and the longer it survives. The smaller the difference between the its ellipse axes lengths and its principal stresses the stabler it is. A standard stress stability measure is what is known as the stress ratio VMi13
@@
Any model for the shape stability can be used and, for illustration purposes, assume a similar form,
@@
The rationale behind this definition is that long narrow cells, for which $f_{c}\to 1$, are very unstable and vice versa. Next, a survivability function is required, whose value determines the configurations occurrence probabilities. It should depend on the stability measures, on the DOFs, and on constraints on the system. Its main aim is to quantify the stability, and thence survivability, of system-wide configurations. The survivability function should take into considerations the observed function collapses and the shape-stress correlations as coupling between $\theta_{c}$ and $\phi_{c}$ at the cell level. In principle, it could also include correlations between the DOFs of nearest-neighbour cells, although searching for such correlations experimentally revealed none Waetal20 ; Suetal25 . Making the survivability function decrease with stability would make it analogous to the free energy in thermal systems. It is natural to use a Boltzmann factor-like exponential to weigh the configurations occurrence probabilities. This is because the stability is an extensive quantity - the overall stability is the sum over the stability measures of the individual cells. The exponential is the only function that accommodates extensivity in that the occurrence probability of a system of independent subsystems is the product of their individual probabilities and its extensive properties are the sum of the individual ones. Finally, the partition function should also be grand-canonical because the number of cells, $N_{c}$, fluctuates. All these specifications are satisfied by
@@
In this expression, $\mu$, $J_{1}$, $J_{2}$, and $J_{3}$ are Lagrange multipliers that reflect the effect of the dynamics-generated noise on, respectively, the number of cells, the cell shape stability, the cell stress stability, and the cell-shape orientational correlations. The choice of a square form for the third term is for simplicity, because it enables analytical calculations, and in the absence of any other data on it. In general, any functional form may be used, as long as it increases with the difference between the angles such that the survivability function decreases monotonically with it. The last term represents the cell-cell interaction, with $\psi_{c^{\prime}c}=\left(\theta_{c}-\theta_{c^{\prime}}\right)^{2}$ the difference between the orientations of neighboring cells $c$ and $c^{\prime}$. More interaction terms between other DOFs of neighbouring cells can be added to $F_{c}$, if necessary.
As the DOFs are continuous, it makes sense to convert the summation in (3) into integrals of ‘densities of states’ over the phase space of DOFs. Denoting the cell DOFs by $\left\{X\right\}=\left\{a,b,\theta,\sigma_{1},\sigma_{2},\phi\right\}$, the densities of states provide the occurrence probabilities of the DOFs and/or the cells stability measures. For the 2D granular systems, some of these distributions are known from modelling and observations. Specifically, $\theta_{c}$ is isotropic on the cell level and $\phi_{c}$ is strongly correlated with it Jietal25 . Thus, these two variable can be replaced by one variable, $\Delta\equiv\theta-\phi$. This variable is normally distributed almost identically around zero for all $k$, $P(\Delta)\sim e^{-\Delta^{2}/\left(2\delta^{2}\right)}$. In principle, the distributions of the stability measures $h$ and $f$ can be obtained from the distributions of the DOFs on which they depend. However, it has been shown that, in these systems, the conditional distributions of $h_{c}$ given $k$, $P_{h}(h|k)$, when scaled by the mean for each cell order, $\bar{h}(k)$, collapse onto a unique Weibull form, $W\left[\hat{h}=h/\bar{h}(k)\right]$ We51 , that is independent of $k$ Jietal25 . A model for this collapse is detailed in the supplemental material SM , predicting almost exactly the experimental observations,
$W(\hat{h})=g_{0}\hat{h}e^{-\frac{g_{0}}{2}\hat{h}^{2}}\ ,$ (5)
with $g_{0}$ a parameter determining transition probabilities between states. This PDF is used in the following calculations. Following the observations in Waetal20 ; Suetal25 and for simplicity, cell-cell correlations are neglected next, $J_{c^{\prime}c}=0$, and the formalism reduces to an ‘independent cells model’. For the illustration here, the conditional density of states of $f(k)$ are assumed to collapse onto a master form of the same form as of $\hat{h}$. An explicit calculation of the partition function, detailed in the supplemental material SM , yields
@@
with the angular brackets denoting averages over all the DOFs. This expression can be used to obtain expectation values of system-wide measurable quantities as functions of the noise parameters and coupling constants in the survivability function,
@@
For example, the expectation values $\langle\hat{f}\rangle$ and $\langle\hat{h}\rangle$ are
$\langle\hat{f}\rangle$ $=\sum\limits_{k=3}^{K}Q_{k}\bar{f}(k)=-\frac{\partial\ln{Z}}{\partial J_{1}}$ (8)
$\langle\hat{h}\rangle$ $=\sum\limits_{k=3}^{K}Q_{k}\bar{h}(k)=-\frac{\partial\ln{Z}}{\partial J_{2}}\ .$ (9)
Explicit forms of these, based on the above assumptions, are derived in detail in the supplemental material SM and are too cumbersome to include here. Also derived in the supplemental material is the (cumbersome) relation
$\langle\hat{f}\rangle=\prod\limits_{k=1}^{K}\left\{\left[\frac{u^{\prime}-v^{\prime}\chi^{\prime}}{u-v\chi}\frac{w-2u\chi+v\chi^{2}}{w^{\prime}-2u^{\prime}\chi^{\prime}+v^{\prime}\chi^{\prime 2}}\right]\right\}^{Q_{k}}\langle\hat{h}\rangle\$, (10)
with all the quantities in this expression defined there. As $\langle\hat{f}\rangle$ and $\langle\hat{h}\rangle$ are system-wide characteristics, (10) is the equivalent of an equation of state. Significantly, this is a relation between properties of the stress and the structure.
The extension of this formalism to three-dimensional (3D) granular SO is conceptually straightforward. As in 2D, cells are the smallest polyhedral voids surrounded by particles in contact. They can be approximated as ellipsoids and their stresses defined as weighted sum of the stresses of their surrounding particles. This would give 10 DOFs per cell (three axes and two angles for each of the shape and stress ellipsoids). Similar ellipsoid and stress stability measures can be defined and $F_{c}$ can be a straightforward 3D version of (4). This, in principle, provides the partition function. However, its evaluation may be more difficult in 3D because of the more complicated cells classification into families. While the 2D cell polygons can be classified only by their order, classification of a 3D polyhedral cell requires more parameters: the number of its faces and the order of each polygonal face. Specifically, a family of cell polyhedra is classified by its $n_{p}$ vertices (contacts), $k_{3}$ triangular faces, $k_{4}$ rectangular faces, and so on up to the highest-order faces, $k_{t}$. Members of the same family have an identical series of such numbers. Nevertheless, this complication is a quantitative, rather than qualitative. It just means that the product over $k$ in (3) needs to be replaced by products over all the cell parameters, but the integrand is the same. This extension is under investigation currently, but it is downstream from the aim here.
@@ IV. The principle and its generalisation
The emergence of SO in granular systems suggests a general principle. SO corresponds to minimising entropy in processes where a minute subset of configurations have a much higher occurrence probabilities than the overwhelmingly larger number of configurations. This interpretation is supported by further numerical and experimental observations in disordered disc assemblies. It was found that the mean area of $k$-cells, when normalised by the area of the regular $k$-polygon of the same edge length, first decreases with $k$ and then increases again MaBl17 ; Suetal20 . The initial decrease is the result of an increasing number of elongated cell configurations, which are stable at low cell orders but become increasingly unstable as $k$ increases. It follows that the effect of stress on the structure and cell stability limits configurational entropy.
This principle holds beyond granular systems. Driving a disordered system generates noise even if thermal fluctuations are negligible. Such systems are also constrained by the environment through boundary conditions and interactions with other dynamic systems, as well as internally by the nature of interactions between their individual components. SO emerges when only a minute subset of system configurations survives the noise and constraints sufficiently long to dominate experimental observations. As the bulk of configurations disintegrate much faster, they are not typically observed. Different driving conditions lead to different noise characteristics and to different SO characteristics, as it is the noise that determines which configurations survive the longest.
V. A general modelling framework
Extending the modelling formalism to more general systems is as follows. (1) define quasi-particles (e.g., the cells); (2) identify the quasi-particles’ DOFs (the cell ellipse axes lengths and orientations, principal stresses and orientation); (3) identify quasi-particle stability measures that depend on the DOFs and whose values decrease with increasing quasi-particle stability (e.g., the cell stresses and shapes); (4) construct a survivability function that decreases monotonically as quasi-particles’ stability increases (e.g., $F_{c}$ in eq. (4)). The value of the survivability function is determined by each system-wide configuration’s stability and it determines the configuration’s occurrence probability and lifetime.